Circular No. 5889

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

4U 0142+614 G. L. Israel, Physics Department, University of Rome; S. Mereghetti,

Istituto di Fisica Cosmica del C.N.R., Milan; and L. Stella, Brera Astronomical Observatory, Milan, communicate: “Using the EXOSAT archival data, we discovered a periodic pulsation from the direction of the two x-ray sources 4U 0142+614 and RX J0146.9+6121. The coherent pulsation, at a heliocentric period of 8.68723 ± 0.00004 s, has been discovered in the Medium Energy (ME) experiment light curves from a 13-hr observation on 1984 Aug. 27. The periodicity is visible (with a 8σ significance) only for energies in the range 1-4 keV (i.e., where most of the x-ray flux is emitted from the optically unidentified source 4U 0142+614). The peak-to-peak amplitude of the modulation is ∼ 7 percent. During the same observation, oscillations at 1455 s in an additional spectral component dominating above 3 keV had been previously reported (White et al. 1987, MNRAS 226, 645). The ME observations of 1985 November-December show a periodicity at 8.666 ± 0.001 s in the range 1-50 keV (Ar + Xe detectors), though at a lower statistical significance. In the 1985 observations, the additional spectral component, probably due to the Be system RX J0146.9+6121 (which lies at 25′ from 4U 0142+614; Mereghetti, Stella, and De Nile 1993, A.Ap. 278, L23) was absent. Thus the pulsations likely originate in the latter source. This is also suggested by ROSAT HRI data obtained on 1991 Feb. 2, in which a periodicity at P = 8.60 s is detected from 4U 0142+614 with a 2σ statistical significance. Further imaging x-ray data are required to confirm that the pulsations are not due to another source within the 90′ field-of-view of the EXOSAT ME. If the 8.7-s periodicity is from 4U 0142+614 and the observed spin-up is confirmed, the lack of an optical counterpart brighter than V ∼ 24 (Steinle et al. 1987, Ap. Space Sci. 131,

687) suggests that this source is a low-mass x-ray binary with an unusually soft spectrum at a distance of several kiloparsecs. The detection of periodic pulsations weakens the phenomenological criterion that an ultra-soft spectral component is a signature of accreting black-hole candidates (White and Marshall 1984, Ap.J. 281, 354).”

COMET MUELLER (1993p) Total visual magnitude estimates by A. Hale, near Cloudcroft, NM

(0.41-m reflector): Oct. 25.41 UT, 12.4; Nov. 6.21, 12.1.

1993 November 12 Daniel W. E. Green